# A Second Main Theorem for Entire Curves Intersecting Three Conics

### v1

@@ Acknowledgements
We also thank Yunling Chen, Alexandre Eremenko, and Yi C. Huang for their suggestions on an earlier version.
The results of this paper were first presented by S.-Y. Xie at The Seminar on Several Complex Variables and Complex Geometry (held from July 27 to August 2, 2025), which was organized by Xiangyu Zhou, Yum-Tong Siu, and John Erik Fornæss.
@@
S.-Y. Xie is partially supported by the National Key R&D Program of China under Grants No. 2023YFA1010500 and No. 2021YFA1003100, as well as by the National Natural Science Foundation of China under Grants No. 12288201 and No. 12471081. D.T. Huynh is supported by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under the grant number 101.04-2025.19
@@ Appendix B A Cohomological Observation
$0\subset S^{m}\,T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})=F^{0}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\subset F^{1}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\subset\cdots\subset F^{\lfloor m/3\rfloor}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\subset E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$
@@
For convenience, here we write $\overline{T}^{*}_{\mathbb{P}^{2}}\coloneqq T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ and $\mathcal{O}(1)\coloneqq\mathcal{O}_{\mathbb{P}^{2}}(1)$.

### v2

@@ Acknowledgements
We also thank Yunling Chen, Alexandre Eremenko, and Yi C. Huang for their suggestions on an earlier version. We thank Steven Lu and Katsutoshi Yamanoi for helpful comments.
The results of this article were first presented by S.-Y. Xie at The Seminar on Several Complex Variables and Complex Geometry (held from July 27 to August 2, 2025), which was organized by Xiangyu Zhou, Yum-Tong Siu, and John Erik Fornæss.
@@
S.-Y. Xie acknowledges partial support from the National Key R&D Program of China under Grants No. 2023YFA1010500 and No. 2021YFA1003100, and from the National Natural Science Foundation of China under Grants No. 12288201 and No. 12471081, as well as support from the Xiaomi Young Talents Program. D.T. Huynh is supported by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under the grant number 101.04-2025.19.
@@ Appendix B A Cohomological Observation
$0\subset S^{m}\,T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})=F^{0}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\subset F^{1}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))\subset\cdots\subset F^{\lfloor m/3\rfloor}(E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C}))=E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$
@@
For convenience, from now on we write $\overline{T}^{*}_{\mathbb{P}^{2}}\coloneqq T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})$ and $\mathcal{O}(1)\coloneqq\mathcal{O}_{\mathbb{P}^{2}}(1)$.